### v4

# Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code

Abstract
Identifying LLM-generated code through watermarking poses a challenge in preserving functional correctness. Previous methods rely on the assumption that watermarking high-entropy tokens effectively maintains output quality. Our analysis reveals a fundamental limitation of this assumption: syntax-critical tokens such as keywords often exhibit the highest entropy, making existing approaches vulnerable to logic corruption. We present STONE, a syntax-aware watermarking method that embeds watermarks only in non-syntactic tokens and preserves code integrity. For rigorous evaluation, we also introduce STEM, a comprehensive metric that balances three critical dimensions: correctness, detectability, and imperceptibility. Across Python, C++, and Java, STONE preserves correctness, sustains strong detectability, and achieves balanced performance with minimal computational overhead. Our implementation is available at https://github.com/inistory/STONE-watermarking.
Marking Code Without Breaking It: Code Watermarking for Detecting LLM-Generated Code
  Jungin Kim   Shinwoo Park   Yo-Sub Han Yonsei University, Seoul, Republic of Korea jungin.kim@yonsei.ac.kr, pshkhh@yonsei.ac.kr, emmous@yonsei.ac.kr
footnotetext: $\star$ Authors equally contributed.footnotetext: $\dagger$ Corresponding author.
1 Introduction
The rapid development of LLMs has significantly improved their ability to generate human-like code  (Zan et al., 2022; Tang et al., 2023; Wei et al., 2023). These advancements have unlocked new possibilities, including software development and automated code generation (Nam et al., 2024; Wang et al., 2024; Guo et al., 2024a). However, this progress has raised new challenges in tracing the provenance of generated code (Yang et al., 2023; Li et al., 2023). As LLM-generated code becomes increasingly indistinguishable from human-written code, researchers have explored watermarking techniques to address this issue. As illustrated in Figure 1, existing approaches often modify syntax-critical tokens, which can introduce syntax errors or alter program behavior. In contrast, STONE preserves these tokens and embeds watermarks only into non-syntactic elements, thereby maintaining structural and functional integrity.
Figure 1: Motivating example of STONE watermarking. Existing methods modify syntax tokens and cause syntax errors, whereas STONE preserves syntax and embeds watermarks without breaking code structure.
LLM watermarking embeds specific patterns into the generated output during the generation process to help identify its origin  (Sander et al., 2024; Golowich and Moitra, 2024; Hu and Huang, 2024) These patterns remain imperceptible to humans yet detectable through algorithmic analysis, thereby promoting greater transparency and accountability in AI-generated content (Guo et al., 2024b; Hou et al., 2024). Among various approaches, the green and red list-based watermarking is the most widely adopted (Kirchenbauer et al., 2023; Zhao et al., 2023; Lee et al., 2024; Lu et al., 2024). These approaches divide token candidates into two lists (green and red) and bias the selection process toward tokens in the green list. Detection involves measuring the proportion of green list tokens in the generated output. While effective for natural language tasks, this approach does not easily apply to code generation, where structural and functional correctness are crucial (Lee et al., 2024; Hoang et al., 2024). Changes to syntax-critical tokens can cause compilation errors or alter program behavior.
We propose STONE (Syntax TOkeN preserving codE watermarking), a method that selectively embeds watermarks in non-syntax tokens. By excluding tokens that are critical to code execution, STONE reduces the risk of functional degradation while embedding robust, detectable patterns. Additionally, we introduce STEM (Summative Test metric for Evaluating code-waterMarking). Existing studies on code watermarking rely on separate evaluation criteria, using different metrics and measurement protocols across methods, hindering fair and consistent comparison of performance. STEM integrates correctness, detectability, and imperceptibility into a unified metric, providing a fair basis for evaluating overall performance while allowing task-specific weighting. We employ STEM to evaluate STONE and compare it with state-of-the-art watermarking methods. Experiments across Python, C++, and Java show that STONE preserves functional correctness, achieves balanced performance across all evaluation criteria, and consistently outperforms prior methods in efficiency.
Figure 2: Entropy values by token type. Tokens that do not fall under keywords, whitespace, types, delimiters, or operators are categorized as etc tokens. The etc category is the primary target for our proposed STONE watermarking method. Token entropy is measured using Qwen2.5-Coder-7B.
2 Motivation and Preliminary Analysis
The recent code watermarking method SWEET (Lee et al., 2024) embeds watermarks in high-entropy tokens. This design is motivated by the intuition that modifying tokens with higher uncertainty is less likely to affect the perceived quality of the generated code. In our preliminary analysis, we examine whether token entropy indeed varies systematically across syntactic categories and how such variation relates to potential impacts on code quality.
We consider two popular benchmarks, MBPP+ and HumanEval+, for Python code generation and completion. Each Python token is classified into one of five syntactic categories (keywords, whitespace, types, delimiters, and operators), with all remaining tokens grouped into an etc category. Details of the categorization procedure are provided in Appendix C.
We compute token entropy for each category using Shannon entropy, following Lee et al. (2024). Given a token sequence $y=(y_{0},y_{1},\ldots,y_{N})$, the entropy at generation step $t$ is defined as:
$H_{t}=-\sum_{i=1}^{|V|}P(y_{t}=v_{i}\mid y_{<t})\log P(y_{t}=v_{i}\mid y_{<t}),$ (1)
where $V$ denotes the vocabulary and $y_{<t}$ the previously generated tokens. This formulation captures the uncertainty of the model’s next-token prediction at each generation step.
Figure 2 shows the mean entropy values for each token category. We observe that keywords exhibit the highest average entropy among all categories. This can be attributed to their diverse roles across control flow, structural definitions, and other syntactic constructs in Python. In contrast, operators, delimiters, types, and whitespace follow more regular and constrained syntactic patterns, resulting in comparatively lower entropy.
Although keywords have high entropy, watermarking them risks altering tokens that are essential to code syntax and semantics. Even small perturbations to such tokens may lead to functional errors or noticeable degradation in code quality. In comparison, the etc category contains tokens that are less critical to core syntactic structure while still exhibiting relatively high entropy. This observation suggests that embedding watermarks in non-syntactic tokens may preserve code quality better than approaches that rely solely on high-entropy tokens.
3 Methodology
3.1 Syntax-Aware Watermarking: STONE
STONE is a code watermarking method designed to preserve the quality of generated code. This section introduces the watermark insertion and detection procedures of STONE.
Inserting Watermarks
We treat keywords, whitespace, types, delimiters, and operators as syntactic elements and embed watermarks only in non-syntactic tokens. Algorithm 1 describes how STONE embeds watermarks during code generation. Given a prompt $\mathbf{x}$ and previously generated tokens $\mathbf{y}_{<t}$, the language model $f_{\text{LM}}$ computes the initial probability distribution $\mathbf{p}_{t}$ at time step $t$. STONE first samples a candidate token $\tilde{y}_{t}$ from the initial probability distribution. If the sampled token is not in the syntax element set, STONE then divides the vocabulary $\mathcal{V}$ into a green and red list according to the green token ratio $\gamma$. It then increases the likelihood of sampling green-list tokens by adding a constant $\delta$ to their logit values. Using the adjusted logits, STONE updates the probability distribution and samples the final token $y_{t}$ for time step $t$.
Algorithm 1 Insertion Algorithm of STONE
1:Prompt $\mathbf{x}$, start token $y_{0}$, last generated tokens $\mathbf{y}_{<t}=(y_{0},\dots,y_{t-1})$, vocabulary set $\mathcal{V}$, language model $f_{\text{LM}}$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, logit adjustment factor $\delta>0$
2:for $t=1,2,3,\dots$ do
3:  Compute the logit vector:
$\mathbf{l}_{t}=f_{\text{LM}}(\mathbf{x},\mathbf{y}_{<t})$
4:  Compute the initial probability vector $\mathbf{p}_{t}$:
$p_{t,i}=\frac{e^{\mathbf{l}_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{\mathbf{l}_{t}[j]}},\quad\text{for }i\in\{1,\dots,|\mathcal{V}|\}$
5:  Sample a candidate token $\tilde{y}_{t}\sim\mathbf{p}_{t}$
6:  if $\tilde{y}_{t}\notin S$ then
7:   Compute a hash of token $y_{t-1}$ as a seed
8:   Split vocabulary indices into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$
9:   Increase logits for indices in $\mathcal{G}_{t}$ by $\delta$:
$\mathbf{l}_{t}[i]\leftarrow\mathbf{l}_{t}[i]+\delta,\quad\text{for }i\in\mathcal{G}_{t}$
10:   Recompute the adjusted probability vector $\mathbf{p}^{\prime}_{t}$:
$p^{\prime}_{t,i}=\frac{e^{\mathbf{l}_{t}[i]}}{\sum_{j=1}^{|\mathcal{V}|}e^{\mathbf{l}_{t}[j]}},\quad\text{for }i\in\{1,\dots,|\mathcal{V}|\}$
11:  else
12:   Set $\mathbf{p}^{\prime}_{t}=\mathbf{p}_{t}$
13:  end if
14:  Sample the final token $y_{t}\sim\mathbf{p}^{\prime}_{t}$
15:end for
Detecting Watermarks
We assess how many green tokens are present in the code, and if this value surpasses a certain threshold, we conclude that the code was generated by an LLM. When identifying potential green tokens, we focus only on non-syntactic elements. Algorithm 2 details the watermark detection process. $N^{E}$ represents the number of etc tokens (non-syntax tokens) in the code, and $N^{E}_{G}$ denotes the number of green tokens among them. We replicate the division of the vocabulary into green and red lists at each time step during code generation to determine whether a specific token is a green token. If the $z$-score exceeds the predefined threshold $z_{\text{threshold}}$, the code is determined to be generated by an LLM.
Algorithm 2 Detection Algorithm of STONE
1:Token sequence $\mathbf{X}=(X_{0},\dots,X_{N-1})$, syntax element set $S$, green list ratio $\gamma\in(0,1)$, $z$-score threshold $z_{\text{threshold}}>0$
2:Initialize $N^{E}\leftarrow 0$, $N^{E}_{G}\leftarrow 0$
3:for $t=1,2,\dots,N-1$ do
4:  if $X_{t}\notin S$ then
5:   $N^{E}\leftarrow N^{E}+1$
6:   Compute a hash of token $X_{t-1}$ as a seed
7:   Split vocabulary indices into $\mathcal{G}_{t}$ and $\mathcal{R}_{t}$
8:   if $X_{t}\in\mathcal{G}_{t}$ then
9:     $N^{E}_{G}\leftarrow N^{E}_{G}+1$
10:   end if
11:  end if
12:end for
13:Compute the $z$-score:
$z=\frac{N^{E}_{G}-\gamma N^{E}}{\sqrt{\gamma(1-\gamma)N^{E}}}$
14:if $z>z_{\text{threshold}}$ then
15:  return True $\triangleright$ Sequence $\mathbf{X}$ is watermarked
16:else
17:  return False $\triangleright$ Sequence $\mathbf{X}$ is not watermarked
18:end if
3.2 Unified Evaluation Metric: STEM
Prior studies on code watermarking typically assess functional correctness, detectability, and imperceptibility independently (Lee et al., 2024; Yang et al., 2024; Li et al., 2024; Guan et al., 2024). They adopt diverse definitions and evaluation protocols, which prevents consistent and fair comparison across methods. For instance, a technique may improve detectability at the cost of correctness, leading to inconsistent evaluations. STEM addresses this issue by unifying the three criteria into a single weighted metric $(\alpha,\beta,\zeta)$, where $\alpha+\beta+\zeta=1$, enabling fair and direct comparison across methods under configurable trade-offs.
$\text{{STEM}}=$ $\;\alpha\cdot\mathrm{Correctness}(C_{wm})$ (2)
$+\beta\cdot\mathrm{Detectability}(C_{wm},C_{H})$
$+\zeta\cdot\mathrm{Imperceptibility}(C_{wm},C).$
Correctness
Correctness is crucial for preventing watermarking from compromising code functionality. Let $C_{wm}$ denote the set of watermarked code samples. We measure functionality preservation using the unbiased estimator of pass@k Chen et al. (2021), which is standard for evaluating code generation correctness:
$\mathrm{Correctness}(C_{wm})=\mathbb{E}_{C_{wm}}\!\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$], (3)
where $n$ corresponds to the total number of generated solutions for a task, and $c$ denotes the subset of those solutions that pass all test cases. This estimator measures the probability that at least one of the $k$ generated solutions is functionally correct. In the context of watermarking, it captures whether watermark insertion introduces execution failures or alters program semantics. A higher correctness score indicates that watermark insertion does not hinder program execution.
Detectability
Detectability measures how effectively a watermark can be identified in generated code. We assess detectability using a $z$-score that quantifies the proportion of green-list tokens, and aggregate scores by computing the AUROC between human-written and watermarked corpora.
For a generated sequence $X_{wm}=(X_{0},X_{1},\dots,X_{N-1})$, the $z$-score of green-token ratio is defined as
$z(X_{wm})=\frac{|X_{wm}|_{G}-\gamma|X_{wm}|}{\sqrt{\gamma(1-\gamma)|X_{wm}|}},$ (4)
where $|X_{wm}|_{G}$ is the number of green tokens in $X_{wm}$ and $\gamma$ is the expected green-token ratio in non-watermarked code. Higher $z$ values indicate stronger evidence that the code is watermarked. The $z$-score serves as a continuous detection statistic, enabling threshold-based discrimination between watermarked and non-watermarked sequences. By varying the threshold $\tau$, different operating points in terms of true and false positive rates are obtained, forming the ROC curve.
Given a corpus of human-written code $C_{H}=\{X^{(i)}_{H}\}$ and watermarked code $C_{wm}=\{X^{(j)}_{wm}\}$, the true positive rate (TPR) and false positive rate (FPR) at threshold $\tau$ are defined as follows:
$\text{TPR}(\tau)$ $=\frac{\sum_{j=1}^{J}\mathbf{1}\!\bigl[z(X_{wm}^{(j)})>\tau\bigr]}{J}.$ (5)
$\text{FPR}(\tau)$ $=\frac{\sum_{i=1}^{I}\mathbf{1}\!\bigl[z(X_{H}^{(i)})>\tau\bigr]}{I}.$ (6)
The area under the ROC curve (AUROC) is then
$\mathrm{Detectability}(C_{wm},C_{H})=\\ \int_{0}^{1}\text{TPR}(\text{FPR})\,d(\text{FPR}),$ (7)
where higher AUROC values indicate stronger separability between watermarked and human-written code.
Imperceptibility
Imperceptibility concerns the preservation of token probability distributions consistent with normal model outputs (Huang and Wan, 2025). Since code LLMs follow stable and highly structured token distributions, distortions introduced by watermarking can form statistical signals that are exploitable by adversaries without access to the secret key. Accordingly, following perplexity-based approaches in text watermarking (Kirchenbauer et al., 2023) and studies on code generation (Magnusson et al., 2024), we define imperceptibility as the deviation in token probability distributions introduced by watermarking, measured through perplexity shifts computed with code LLMs.
Let $C_{wm}$ and $C$ denote the corpora of watermarked and non-watermarked code, respectively. Perplexity is computed over the watermarked corpus ${C}_{wm}=\{X_{wm}^{(1)},\dots,X_{wm}^{(J)}\}$ as:
$\operatorname{PPL}(C_{wm})=\\ \frac{1}{\lvert C_{wm}\rvert}\sum_{j=1}^{\lvert C_{wm}\rvert}\exp\!\Bigl(-\mathbb{E}_{i}\!\left[\log P\!\left(y_{i}^{(j)}\mid y_{<i}^{(j)}\right)\right]\Bigr),$ (8)
where $P(y_{i}^{(j)}|y_{<i}^{(j)})$ represents the probability of token $y_{i}^{(j)}$ given its preceding context. An effectively concealed watermark minimizes perplexity shifts, ensuring imperceptibility. We define the imperceptibility metric as:
$\mathrm{Imperceptibility}(C_{wm},C)=\\ 1-\frac{\bigl|\operatorname{PPL}(C_{wm})-\operatorname{PPL}(C)\bigr|}{\operatorname{PPL}(C)}.$ (9)
This metric quantifies the perplexity-based difference between LLM-generated code with watermark ($C_{wm}$) and non-watermarked code ($C$). It computes the relative difference by normalizing the absolute gap by the perplexity of $C$, providing a scale-invariant measure of distributional shift. Higher values indicate minimal impact on token distributions and thus stronger imperceptibility, while lower or negative values reflect greater disruption due to watermark insertion.
4 Experimental Setup
Datasets.
For our experiments, we evaluate our approach across three programming languages: Python, C++, and Java. We use four benchmark datasets: HumanEval+ and MBPP+ (Liu et al., 2023) for Python, and HumanEvalPack (Muennighoff et al., 2024) for C++ and Java. Table 1 reports token-length statistics of the evaluation datasets, based on the Qwen2.5-Coder-7B tokenizer. The datasets span from short snippets to long sequences, allowing evaluation under varying code lengths and structural complexity.
Dataset Problems Test Cases Max Min Mean Std.
MBPP+ 399 105 341 11 40.25 34.91
HumanEval+ 164 764 558 43 188.28 87.32
HEP-C++ 164 764 697 42 223.10 107.57
HEP-Java 164 764 660 56 237.36 105.09
Table 1: Token length statistics of solution code across evaluation datasets, along with the total number of problems and the average number of test cases per problem.
Baselines.
We consider the following three watermarking approaches for comparison. (1) KGW (Kirchenbauer et al., 2023) generates text by dividing the vocabulary into a green and red list at each time step of token generation, increasing the probability of generating tokens from the green list. (2) EWD (Lu et al., 2024) leverages entropy-based token weighting during detection, giving more influence to high-entropy tokens, thus enhancing detection in texts with varying entropy distributions. (3) SWEET (Lee et al., 2024) selectively embeds watermarks in high-entropy tokens, addressing detection challenges posed by the low entropy in code. We exclude CodeIP (Guan et al., 2024) from our baselines as it requires additional training for watermark insertion, while our approach and the selected baselines are training-free methods. As part of our comparative analysis, we reproduce CodeIP and report its experimental results in Appendix A.
Base Model.
We use Qwen2.5-Coder-7B (Hui et al., 2024) for our experiments. We also report additional results with Llama-3.1-8B in Appendix B, which further demonstrate the generalizability of our approach across different model architectures.
Evaluation Metric.
We evaluate watermarking methods across three dimensions: Correctness, Detectability, and Imperceptibility. Correctness uses pass@k (k=1,5), detectability is measured by AUROC with a $z$-score test, and imperceptibility by perplexity computed with StarCoder2-7B. We use the STEM metric to integrate these three axes into a single composite score. For the main results (Tables 2 and 3), we adopt an equal-weight configuration $(\tfrac{1}{3},\tfrac{1}{3},\tfrac{1}{3})$. We further analyze weight settings with alternative configurations in Appendix D.
Implementation Details.
During code generation, we apply top-$k$ sampling with $k=50$, restricting the candidate token pool to the top 50 tokens with the highest logits and normalizing their probabilities before sampling. The temperature, which controls the sharpness of the token distribution, is fixed at 1.0. We set $\gamma=0.5$ and $\delta=1.0$ for the MBPP+ dataset, while for the HumanEval+ and HumanEvalPack datasets (C++ and Java), we set $\gamma=0.5$ and $\delta=0.5$. We conduct all experiments on an NVIDIA A6000 GPU.
Method MBPP+ HumanEval+
Correctness Detectability Imperceptibility STEM Correctness Detectability Imperceptibility STEM
KGW 0.499 0.831 0.994 0.775 0.573 0.523 0.986 0.694
EWD 0.499 0.965 0.994 0.819 0.573 0.730 0.986 0.763
SWEET 0.502 0.867 0.992 0.787 0.574 0.710 0.978 0.754
STONE 0.571 0.982 0.990 0.848 0.587 0.777 0.978 0.781
Table 2: Experimental results on MBPP+ and HumanEval+. Metrics include correctness, detectability, imperceptibility, and the integrated STEM score under the equal-weight configuration.
Method HEP-C++ HEP-Java
Correctness Detectability Imperceptibility STEM Correctness Detectability Imperceptibility STEM
KGW 0.576 0.621 0.993 0.730 0.387 0.546 0.993 0.642
EWD 0.576 0.681 0.993 0.750 0.387 0.646 0.993 0.675
SWEET 0.584 0.641 0.979 0.735 0.413 0.580 0.901 0.631
STONE 0.622 0.729 0.990 0.780 0.445 0.721 0.979 0.715
Table 3: Experimental results on HumanEvalPack-C++ and HumanEvalPack-Java. Metrics include correctness, detectability, imperceptibility, and the integrated STEM score under the equal-weight configuration.
5 Experiments and Results
We pose three research questions and provide a thorough analysis of each:
  • RQ1: Can STONE preserve functional correctness?
  • RQ2: Does STONE achieve balanced performance across correctness, detectability, and imperceptibility as measured by STEM?
  • RQ3: How efficient is STONE in watermark insertion and detection?
We address these questions through three analytical dimensions: (1) Functionality Preservation, which verifies that watermark insertion does not alter program behavior; (2) Balanced Performance under STEM, which examines whether STONE maintains a well-balanced trade-off among correctness, detectability, and imperceptibility; and (3) Efficiency, which measures the computational overhead of watermark insertion and detection.
RQ1
We examine whether STONE embeds watermarks without affecting the functional correctness of generated code. Tables 2 and 3 show that STONE improves correctness by 7.57% over SWEET on average across all benchmarks. The token-level analysis in Appendix E indicates that the entropy-based token selection method used by SWEET often chooses syntax-related elements—such as delimiters, whitespace, keywords, types, and operators—that are essential for correct execution. For example, modifying delimiters like colons (‘:’) or brackets (‘’, ‘[]’) can cause parsing errors, and changing operators or keywords (e.g., replacing ‘+’ with ‘-’ or ‘True’ with ‘False’) can alter program logic or control flow. Even under the optimal configuration, syntax tokens account for 12.6% of the tokens selected for watermarking in the HumanEval+ dataset, with delimiters and whitespace forming the majority. This inclusion of syntax elements explains the observed reduction in correctness for SWEET. In contrast, STONE excludes syntax tokens and embeds watermarks only into non-syntactic elements, which prevents structural and logical errors during execution and allows the generated code to preserve its functional behavior.
RQ2
We examine whether STONE achieves balanced performance across correctness, detectability, and imperceptibility. Baselines such as KGW and EWD watermark every token, which dilutes the watermark signal and lowers AUROC scores. SWEET modifies high-entropy tokens, including syntax elements, improving detectability but reducing syntactic stability. In contrast, STONE embeds watermarks only in non-syntax tokens, preserving syntactic integrity while keeping the signal distinct. This selective design enables STONE to achieve higher detection accuracy without compromising code imperceptibility.
Imperceptibility results reinforce this finding. KGW and EWD maintain reasonable imperceptibility by uniformly adjusting token probabilities, but their adjustments can still introduce subtle artifacts. SWEET often produces more noticeable distributional shifts because it alters syntax tokens. STONE, in contrast, maintains low perplexity differences by leaving syntax untouched and modifying only non-syntax tokens. As a result, STONE maintains imperceptibility without sacrificing correctness or detection robustness. We quantify this balanced behavior using the STEM metric, which under the equal-weight configuration provides a neutral evaluation across correctness, detectability, and imperceptibility. Unlike other watermarking methods that exhibit clear trade-offs among these criteria, STONE achieves consistently strong overall performance. A comprehensive grid analysis across 66 weight settings (Appendix D) further confirms that this advantage remains stable and does not depend on any specific weighting scheme.
Method MBPP+ HumanEval+ HEP-C++ HEP-Java
Insertion Detection Insertion Detection Insertion Detection Insertion Detection
KGW 3320 12.90 1268 4.19 1308 8.01 1506 8.95
EWD 3320 100.43 1268 34.51 1308 56.85 1506 59.84
SWEET 3300 100.94 1270 34.68 1454 58.08 1480 59.22
STONE 3266 13.27 1277 4.62 1300 8.48 1459 9.24
Table 4: Comparison of total insertion and detection time (measured in seconds) across different watermarking methods. The reported measurement is performed at the dataset level.
RQ3
We analyze the computational overhead of watermarking in terms of insertion and detection time. Insertion time denotes the duration required to generate watermarked code, whereas detection time measures the time needed to verify a watermark. Table 4 summarizes the results for each method. All approaches exhibit comparable insertion times, but STONE achieves detection speeds that are on average 86% faster than those of SWEET and EWD at the dataset level. This improvement results from differences in detection mechanisms. SWEET and EWD employ entropy-based approaches that compute token probability distributions, which substantially increase computational overhead. In contrast, STONE verifies watermarks using a predefined green list of non-syntax tokens, reducing detection time while maintaining reliable detectability.
6 Analysis
6.1 Trade-off Between Code Quality and Detection Performance
We analyze the trade-off between code quality and watermark detectability as influenced by two watermarking parameters. The parameter $\gamma$ controls the proportion of the vocabulary designated as the green list, while $\delta$ determines the extent to which the selection probability of green tokens is amplified. We conduct an analysis on the HumanEval+ dataset by varying the values of $\delta$ and $\gamma$. Figure 3 shows the results in a bubble chart, revealing a clear trade-off between code quality and detectability. The circular markers represent the pass@1 and detection performance (AUROC) of STONE, while the triangular markers indicate those of SWEET. Among the four regions divided by the two black dashed lines, the upper-right quadrant corresponds to configurations where both code quality and detection performance are high, indicating a desirable balance between utility and detectability. We observe that, while only a single configuration of SWEET falls into this balanced region, multiple configurations of STONE—spanning various combinations of $\gamma$ and $\delta$—lie within it. This indicates that STONE achieves a more balanced trade-off between functionality and watermark detectability compared to SWEET.
Figure 3: The trade-off between pass@1 (Y-axis) and detectability (X-axis). We analyze the impact of the green list ratio $\gamma$ (indicated by marker color) and the watermark strength $\delta$ (indicated by marker size). Circular and triangular markers represent STONE and SWEET, respectively.
6.2 Robustness under Adversarial Attacks
Although a watermark is embedded in the code, a malicious user may remove it by modifying the code. Therefore, the watermark must remain robust against attacks like code modifications. We assess the resilience of watermarks of STONE against code alterations by applying two types of attacks: (1) a code refactoring attack using a commercial service [^1] https://codepal.ai/code-refactor and (2) a code paraphrasing attack using ChatGPT (GPT-4o) [^2] Prompt: Please paraphrase the following code.. We conduct this analysis on the HumanEval+ dataset. Figure 4 shows the classification performance (i.e., detectability) in distinguishing between watermarked LLM-generated code and human-written code after these attacks are applied to code watermarked by STONE and SWEET. The experimental results show that STONE is more robust than SWEET against both types of attacks. Refactoring reorganizes code while keeping syntax boundaries, so STONE preserves many watermarks inserted into non-syntax tokens. Paraphrasing changes variable names and expressions, which directly correspond to the embedding targets of STONE, and thus decreases detectability. SWEET shows the opposite tendency because refactoring alters its watermarking targets. STONE maintains higher robustness under both attacks by embedding watermarks into a syntax-filtered token space that remains stable under structural and lexical modifications. This result confirms that STONE maintains consistent detection robustness across different attack types.
Figure 4: We evaluate robustness by applying code refactoring and paraphrasing attacks to watermarked code and measuring detection performance. Results compare the resilience of SWEET and STONE under both attacks.
7 Related Work
Post-hoc detection methods (Xu and Sheng, 2024; Shi et al., 2024) identify machine-generated code after generation using statistical cues such as perplexity or syntax differences. In contrast, watermarking embeds provenance signals during generation to enable direct identification (Kirchenbauer et al., 2023; Guan et al., 2024; Park et al., 2025). Our work focuses on the latter approach, developing a syntax-aware watermarking method for code generation.
Among watermarking approaches, Lee et al. (2024) proposed SWEET, a method that selectively embeds watermarks in high-entropy tokens. While SWEET demonstrated improvements in detection capability and code quality, our preliminary analysis revealed that approximately 12.6% of these high-entropy tokens correspond to syntax elements such as reserved keywords. Modifying these critical tokens can disrupt program logic, resulting in degraded functional correctness. Guan et al. (2024) introduced CodeIP, a grammar-guided watermarking approach by utilizing type prediction to maintain syntactic validity. This method ensured that the generated code remained grammatically correct. However, the reliance on type prediction requires significant computational resources, limiting its practicality for large-scale code generation tasks. In contrast to these approaches, our method explicitly excludes syntax-critical tokens from watermark insertion. Our approach ensures that essential structural elements, like keywords and operators, remain unchanged, thereby preserving functional correctness without additional computational overhead. ACW (Li et al., 2024) and SrcMarker (Yang et al., 2024) employed watermarking techniques based on code transformations, such as semantic-preserving edits and variable renaming. While these approaches preserved program functionality, the resulting modifications often introduce unnatural code patterns, which can increase the risk of watermark removal by adversaries.
8 Conclusion
We present STONE, a syntax-aware watermarking method that preserves functional correctness by excluding syntax-critical tokens during watermark insertion. We also propose STEM, a unified evaluation metric integrating correctness, detectability, and imperceptibility for balanced assessment. Experiments on Python, C++, and Java show that STONE outperforms baseline methods without degrading code quality and reduces watermark detection time compared with entropy-based approaches. Evaluation under the STEM metric demonstrates balanced performance across weight configurations, and further analysis confirms that STONE maintains stable and robust performance under code modifications.
Limitations
While STONE provides a robust approach to functionality-preserving watermarking, it has several limitations.
First, the watermarking capacity is based on the density of non-syntactic tokens. The conventional coding styles in our benchmark datasets provided ample opportunities for watermark insertion. This efficacy, however, can be compromised in unconventional code, such as obfuscated scripts or solutions for code golf, where such tokens are sparse. Under these conditions, the embedded watermark may prove too sparse for reliable detection.
Second, although our evaluation demonstrates resilience to general attacks like paraphrasing and refactoring, a more sophisticated threat model considers an adversary with specific knowledge of the STONE algorithm. Such an adversary may directly attack non-syntactic tokens by systematically renaming all variables and removing comments. This class of attack poses a more significant threat to watermark integrity than general modifications. Developing countermeasures against such targeted, algorithm-aware attacks remains an important direction for future research.
Acknowledgments
This research was supported by the NRF grant (RS-2025-00562134) and the AI Graduate School Program (RS-2020-II201361) funded by the Korean government.
Appendix A Performance Analysis of CodeIP
[…]
Table 5: Evaluation of CodeIP on four datasets. Correctness: pass@1; Detectability: AUROC; Imperceptibility: PPL change from non-watermarked to watermarked code.
Table 5 reports the reproduced results for CodeIP, a grammar-guided multi-bit watermarking method that employs a type prediction model trained on CodeSearchNet for grammar-aware token selection during code generation. All experiments are based on the original implementation and evaluated on four benchmarks: MBPP+, HumanEval+, HumanEvalPack-C++, and HumanEvalPack-Java.
CodeIP maintains high detectability, achieving AUROC values above 0.94 across all datasets. This result indicates that grammar-based watermark insertion enables reliable message extraction. However, this detection capability is accompanied by substantial drops in correctness and imperceptibility. The pass@1 scores decrease sharply, approaching zero on HumanEval+ and HumanEvalPack-C++, which shows that syntactic validity enforced by the type predictor does not necessarily imply functional correctness. The large perplexity gap of up to +4.9 indicates that the interaction between watermarking and type-prediction constraints can shift the token distribution, leading to deviations from the model’s natural generation pattern.
Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ HEP-C++ HEP-Java
$(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.270 0.296 0.122 0.018
(0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.241 0.274 0.095 0.041
(0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.381 0.361 0.267 0.145
(0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.325 0.331 0.174 0.084
Table 6: STEM scores of CodeIP under different $(\alpha,\beta,\zeta)$ on MBPP+, HumanEval+, HumanEvalPack-C++, and HumanEvalPack-Java.
Table 6 presents the STEM scores of CodeIP under different weight settings. Across all configurations, detectability remains consistently high, while correctness- or imperceptibility-focused STEM scores are significantly lower than those of other methods. Even under weight settings emphasizing correctness or imperceptibility, CodeIP does not achieve balanced performance, showing limited robustness under varying evaluation preferences. These findings demonstrate that maximizing watermark detectability alone severely degrades code quality. In contrast, STONE embeds watermarks only within non-syntax tokens and achieves stable performance across correctness, detectability, and imperceptibility as captured by the STEM metric.
Appendix B Experimental Results on Llama-3.1-8B
Tables 7 and 8 summarize the experimental results obtained with the Llama-3.1-8B model on MBPP+ and HumanEval+. STONE achieves the highest STEM scores across all datasets and weight settings, demonstrating that the non-syntax watermarking strategy generalizes well to larger models. EWD shows slightly higher detectability on HumanEval+, which may be related to differences in the output distribution of Llama-3.1-8B under uniform insertion. In this setting, the cumulative bias introduced by EWD is more effectively reinforced across long and diverse sequences, leading to a small gain in detectability. Such an effect is not observed under Qwen2.5-Coder-7B, whose flatter distributions yield comparable results between EWD and SWEET. Overall, STONE preserves stronger correctness and achieves a more balanced trade-off, resulting in the highest overall STEM scores.
Across different weight configurations, STONE maintains stable performance even when correctness or imperceptibility is prioritized. The method remains competitive under detectability-oriented settings without functional degradation. This consistency confirms that syntax-aware watermarking effectively preserves code quality while sustaining robust watermark signals.
Method MBPP+ HumanEval+ Correct. Detect. Imperc. Correct. Detect. Imperc. KGW 0.414 0.831 0.953 0.323 0.507 0.989 EWD 0.414 0.931 0.953 0.323 0.755 0.989 SWEET 0.404 0.927 0.968 0.354 0.711 0.978 STONE 0.446 0.945 0.954 0.372 0.741 0.983
Table 7: Experimental results on MBPP+ and HumanEval+ (Llama-3.1-8B): Correctness, Detectability, and Imperceptibility.
  Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ KGW EWD SWEET STONE KGW EWD SWEET STONE $(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.733 0.766 0.766 0.782 0.606 0.689 0.681 0.699 (0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.653 0.678 0.676 0.698 0.535 0.598 0.599 0.617 (0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.757 0.807 0.806 0.823 0.582 0.706 0.688 0.709 (0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.788 0.813 0.817 0.825 0.702 0.764 0.755 0.770
Table 8: STEM scores on MBPP+ and HumanEval+ (Llama-3.1-8B) under different weight settings $(\alpha,\beta,\zeta)$. The left column lists weights for correctness, detectability, and imperceptibility.
Category Python C++ Java
Keywords True, False, None, and, as, assert, async, await, break, class, continue, def, del, elif, else, except, finally, for, from, global, if, import, in, is, lambda, nonlocal, not, or, pass, raise, return, try, while, with, yield alignas, alignof, and, and_eq, asm, auto, bitand, bitor, break, case, catch, class, compl, concept, const, consteval, constexpr, constinit, const_cast, continue, co_await, co_return, co_yield, decltype, default, delete, do, dynamic_cast, else, enum, explicit, export, extern, false, for, friend, goto, if, inline, mutable, namespace, new, noexcept, not, not_eq, nullptr, operator, or, or_eq, private, protected, public, register, reinterpret_cast, requires, return, sizeof, static, static_assert, static_cast, struct, switch, template, this, thread_local, throw, true, try, typedef, typeid, typename, union, using, virtual, volatile, while, xor, xor_eq, override abstract, assert, break, case, catch, class, const, continue, default, do, else, enum, extends, final, finally, for, goto, if, implements, import, instanceof, interface, native, new, null, package, private, protected, public, return, static, strictfp, super, switch, synchronized, this, throw, throws, transient, try, void, volatile, while, true, false
[…]
Types int, float, complex, str, bytes, bool, list, tuple, set, dict, NoneType int, float, double, bool, char, short, long, void, unsigned, signed, size_t, ptrdiff_t, wchar_t, char8_t, char16_t, char32_t byte, short, int, long, float, double, boolean, char, String, Object
Delimiters (, ), [, ], {, }, ‘,’, :, ‘.’, ;, @, ->, … (, ), [, ], {, }, ‘,’, :, ‘.’, ;, ->, ::, … (, ), [, ], {, }, ‘,’, :, ‘.’, ;, @, ->, ::, …
[…]
Table 9: Comparison of five categories of syntax elements across Python, C++, and Java.
  Weight $(\alpha,\beta,\zeta)$ MBPP+ HumanEval+ HEP-C++ HEP-Java KGW EWD SWEET STONE KGW EWD SWEET STONE KGW EWD SWEET STONE KGW EWD SWEET STONE $(\tfrac{1}{3},\,\tfrac{1}{3},\,\tfrac{1}{3})$ 0.775 0.819 0.787 0.848 0.694 0.763 0.754 0.781 0.730 0.750 0.735 0.780 0.642 0.675 0.631 0.715 (0.5, 0.25, 0.25)(0.5,\,0.25,\,0.25) 0.706 0.739 0.716 0.778 0.664 0.716 0.709 0.732 0.692 0.706 0.697 0.741 0.578 0.603 0.577 0.648 (0.25, 0.5, 0.25)(0.25,\,0.5,\,0.25) 0.789 0.856 0.807 0.881 0.651 0.755 0.743 0.780 0.703 0.733 0.711 0.768 0.618 0.668 0.618 0.716 (0.25, 0.25, 0.5)(0.25,\,0.25,\,0.5) 0.830 0.863 0.838 0.883 0.767 0.810 0.808 0.830 0.796 0.811 0.796 0.833 0.730 0.755 0.699 0.781
Table 10: STEM scores under different weight settings $(\alpha,\beta,\zeta)$. The left column lists weights for correctness, detectability, and imperceptibility. Bold indicates the best score in each row.
Appendix C Syntax Elements in Python, C++, and Java
Table 9 categorizes the syntactic elements into five groups: keywords, whitespace, types, delimiters, and operators. The selection of syntactic elements reflects the distinct characteristics of each programming language.
Entropy Threshold MBPP+ HumanEval+ HEP-C++ HEP-Java
Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%) Selected Tokens (%) Syntax Tokens (%)
0.7 26.03 12.82 35.28 12.84 30.10 12.99 26.60 11.98
0.8 21.40 12.59 32.67 12.24 25.93 12.35 22.47 11.06
0.9 19.64 11.39 28.98 12.60 22.65 11.79 18.52 11.68
1.0 16.59 11.39 24.84 11.83 18.96 10.56 16.13 11.15
1.1 14.94 10.30 23.22 11.81 15.71 9.82 14.39 10.91
Table 11: Token selection statistics of SWEET across four datasets. Selected Tokens (%) denotes the proportion of tokens exceeding the entropy threshold out of all tokens, and Syntax Tokens (%) represents the percentage of syntax-related tokens among those selected. Bold values indicate the optimal entropy threshold used in SWEET.
Appendix D Comprehensive Evaluation across STEM Weight Configurations
Table 10 reports representative STEM variants using equal weights and three dimension-focused settings. We further explore all valid weight configurations where $\alpha,\beta,\zeta\in\{0.0,0.1,\dots,1.0\}$ and $\alpha+\beta+\zeta=1$, resulting in $66$ combinations. For each configuration, we compute the weighted STEM score and record the method with the highest value. Across all datasets, STONE performs consistently well throughout the entire weight space, ranking first in $97.0\%$ of configurations on MBPP+, $90.9\%$ on HumanEval+, $98.5\%$ on HumanEvalPack-C++, and $95.5\%$ on HumanEvalPack-Java. This consistency indicates that the advantage of STONE does not depend on specific weight choices but generalizes across diverse evaluation preferences, demonstrating robust performance under varying priorities of correctness, detectability, and imperceptibility.
Appendix E Syntax Token Coverage in SWEET
Table 11 presents an analysis of token selection by SWEET using two measures: (1) the proportion of selected tokens exceeding the entropy threshold among all generated tokens, and (2) the proportion of syntax-related tokens among those selected. We use the optimal settings ($\gamma=0.25$, $\delta=3.0$, entropy threshold =0.9=0.9) and evaluate threshold values in the range $\{0.7,0.8,0.9,1.0,1.1\}$. The results show an inverse relationship between the threshold and token coverage: lower thresholds select more tokens overall but include more syntax elements, whereas higher thresholds reduce syntax token inclusion. This pattern shows that SWEET effectively targets high-entropy regions but does not completely avoid syntax-related tokens.
Even under the optimal configuration, a notable portion of syntax tokens remains among those selected for watermarking. In the HumanEval+ dataset, syntax tokens account for 12.6% of all selected tokens, including delimiters (49.29%), whitespace (38.17%), keywords (9.44%), types (3.17%), and operators (2.78%). This indicates that delimiters and whitespace dominate syntax-related selections, which directly affects code execution.
Category Examples
Keywords def, or, import, for, True, assert, return, is, pass, None, False, in, not, if, from
[…]
Types int, set, str
Delimiters . , : , {, }, [, ], (, )
Operators *, %, =, -, **, &, ^, /, +, |
Table 12: Examples of syntax tokens selected for watermarking in SWEET.
Table 12 lists representative examples. Delimiters such as colons (‘:’) and commas (‘,’), essential for code structure, can cause parsing errors when modified (e.g., omitting a colon in ‘for i in range(10):’ leads to a failure). Although keywords and operators appear less frequently, their modification can still alter program behavior, for example, replacing ‘True’ with ‘False’ or ‘+’ with ‘-’. Minimizing the inclusion of such syntax-critical tokens can make syntax-aware watermarking more reliable and less disruptive to functionality.
Appendix F Comparison with Non-watermarked Code
We evaluate the functional impact of watermarking by comparing the correctness of code generated with and without watermark insertion. Non-watermarked code provides an empirical upper bound in our evaluation setting. STONE achieves scores nearly identical to this bound, showing that watermark insertion has minimal influence on code execution. Among all watermarking methods, STONE reduces correctness the least, which indicates that it embeds provenance information effectively while preserving the original program behavior.
Method MBPP+ HumanEval+
No watermark 0.571 0.595
KGW 0.499 0.573
EWD 0.499 0.573
SWEET 0.502 0.574
STONE 0.571 0.587
Table 13: Correctness comparison between non-watermarked and watermarked code. Correctness is measured as pass@1 on MBPP+ and HumanEval+.